# A Matlab-based Toolbox for Automatic EMT Modeling and Small-Signal Stability Analysis of Modern Power Systems

### v1

@@ IV-D User-Defined Elements
Despite STAMP including an extensive library of pre-defined models, not all possible power system devices and control structures are modeled. For this reason, a feature to seamlessly incorporate user-defined elements is also considered. Moreover, sometimes the differential equations of the components are not available due to confidentiality reasons. In this case, the state-space models can be provided in numerical form in order to conceal confidential information, or directly estimated from other types of black-box models [46, 47].
@@ V-B Power Flow Calculation
The operations of STAMP start by solving a power flow problem, whose solution is subsequently used in two ways, as illustrated in Fig. 1. Firstly, it is used for the calculation of the equilibrium point around which the system is linearised. Secondly, it is used to initialize the states of the non-linear system, thus avoiding computationally inefficient TDSs until the model reaches a steady state [48].
If the desired power flow solution is already available, it can be provided directly as input data. Alternatively, STAMP offers several integrated options for the power flow calculation. Aside from individually developed routines, an integrated interface with established, open-source software that is tailored for power flow calculations is provided. Specifically, the user can find designated data parsers and interfaces with MATPOWER and MATACDC for traditional, and AC/DC hybrid power flow calculations, respectively [49, 50].
@@ V-D Non-linear TDS
  • Non-linear model initialization: STAMP includes a feature of automatically initializing the non-linear EMT models. The power flow solution is the input for the initialization algorithm with all the initialization variables being computed using predefined Matlab-based functions for each element included in the library [51].
@@ VI-A WSCC System
In order to illustrate the comprehensive stability analysis capabilities of STAMP, a case study based on the well-known WSCC system with increased penetration of power converters was performed. Compared with the original system, two of the standard SGs were substituted by two VSCs, one operating in droop-based, GFM mode, and one operating in GFL mode. Their control structures and modeling equations are provided in [28, 52], respectively. The remaining SG was modeled with a standard 6th-order machine model and was equipped with an AC4A exciter and an IEEEG1 turbine governor [5, 24]. The above generation mix is representative of modern power systems and introduces various dynamic phenomena of different timescales that can be analyzed using STAMP. Fig. 4 shows the single-line diagram of the system under study.
@@
STAMP offers the possibility of performing parametric stability analysis in the frequency domain. Fig. 7 shows the various implemented options to depict the converter operation in the frequency domain, namely the converter admittance expressed in Park and positive sequence coordinates (in Fig. 7(a) and Fig. 7(b), respectively), as well as the converter passivity index (in Fig. 7(c)). The results of the frequency domain analysis are consistent with the eigenvalue analysis, demonstrating that the admittance angle of the converter is outside of the (-90°$\degree$,90°$\degree$) range in both coordinate systems and for the lowest value of the controller time constant. Additionally, the frequency region where this occurs coincides with the frequency of the unstable modes of Fig. 6(c). Finally, the passivity plot of Fig. 7(c) shows that the converter is non-passive in the same region, implying adverse interactions with the rest of the network in this frequency range and potential instability. The above frequency domain analysis features allow the seamless incorporation of black-box models to the stability analysis, provided that their respective frequency response is given to the software as an input, as explained in Section IV-D.
@@ VI-B INELFE System
This case study aims to showcase the capability of STAMP to perform analysis on AC/DC hybrid power systems and is based on the real-world HVDC interconnection between Spain and France, termed the INELFE link [53]. During the $16^{\text{th}}$ of June 2015, the link suffered a tripping caused from a harmonic instability event, which this case study aims to reproduce [54, 55].
Fig. 8 shows the schematic of the system under study. It is comprised of an HVDC link, two IPCs, rated at 1000 MW, 400 kV each, a double transmission line, an inductive load, and two Thevenin circuits, representing the two interconnected AC networks. IPC-1 regulates the DC voltage at the HVDC link while IPC-2 controls the active and reactive power interchanged with the AC grid. The two transmission lines are modeled as $\pi$-section circuits and the two Thevenin circuits have the same impedance parameters, namely $R_{th}=0.001$ pu and $X_{th}=0.01$ pu. The current loop for both IPCs is tuned with the internal model control method with a time constant of 1 ms [56, 57]. Both IPCs include a control delay of 50 $\mu$s and a zero-order hold of 10 $\mu$s, modeled as in [35].
@@ VI-C Case Study Computation Time
Power flow calculation 0.081 s 0.167 s 0.216 s 0.334 s
Subsystems linear model calculation 0.441 s 0.277 s 2.044 s 6.242 s
Linear model interconnection 0.009 s 0.008 s 0.36 s 2.528 s
Eigenvalue calculation 0.1 s 0.109 s 0.616 s 3.895 s
@@
In this Section, the computation time that STAMP requires its operations is presented. These operations include the pre-calculation of the power flow for the linearization point evaluation, the calculation of the state space matrices of each individual linear subsystem, their interconnection and the final calculation of the eigenvalues from the complete system state matrix. The selected systems include the ones of Section VI-A and VI-B, as well as of two systems of larger size in order to demonstrate the scalability of the toolbox. The two larger models are a variation of the well-known New England benchmark, where 35% of the SGs are substituted with IBRs, as well as a reduced version of the Midwest benchmark system. Both the static and dynamic data that were used for the calculations can be found in [22]. All calculations were performed in a desktop computer, equipped with an AMD Ryzen Threadripper 2950X 16-core processor, operating at 3.5 GHz and a RAM of 32 GB.

### v2

@@ IV-D User-Defined Elements
Despite STAMP including an extensive library of pre-defined models, not all possible power system devices and control structures are modeled. For this reason, a feature to seamlessly incorporate user-defined elements is also considered. Moreover, sometimes the differential equations of the components are not available due to confidentiality reasons. In this case, the state-space models can be provided in numerical form in order to conceal confidential information, or directly estimated from other types of black-box models [smith2024black, cifuentes2021black].
@@ V-B Power Flow Calculation
The operations of STAMP start by solving a power flow problem, whose solution is subsequently used in two ways, as illustrated in Fig. 1. Firstly, it is used for the calculation of the equilibrium point around which the system is linearised. Secondly, it is used to initialize the states of the non-linear system, thus avoiding computationally inefficient TDSs until the model reaches a steady state [alican2023initialization].
If the desired power flow solution is already available, it can be provided directly as input data. Alternatively, STAMP offers several integrated options for the power flow calculation. Aside from individually developed routines, an integrated interface with established, open-source software that is tailored for power flow calculations is provided. Specifically, the user can find designated data parsers and interfaces with MATPOWER and MATACDC for traditional, and AC/DC hybrid power flow calculations, respectively [matpower, matacdc].
@@ V-D Non-linear TDS
  • Non-linear model initialization: STAMP includes a feature of automatically initializing the non-linear EMT models. The power flow solution is the input for the initialization algorithm with all the initialization variables being computed using predefined Matlab-based functions for each element included in the library [ding2024holistic].
@@ VI-A WSCC System
In order to illustrate the comprehensive stability analysis capabilities of STAMP, a case study based on the well-known WSCC system with increased penetration of power converters was performed. Compared with the original system, two of the standard SGs were substituted by two VSCs, one operating in droop-based, GFM mode, and one operating in GFL mode. Their control structures and modeling equations are provided in [pogaku2007modeling, collados], respectively. The remaining SG was modeled with a standard 6th-order machine model and was equipped with an AC4A exciter and an IEEEG1 turbine governor [kundur, milano2010power]. The above generation mix is representative of modern power systems and introduces various dynamic phenomena of different timescales that can be analyzed using STAMP. Fig. 4 shows the single-line diagram of the system under study.
@@
STAMP offers the possibility of performing parametric stability analysis in the frequency domain. Fig. 7 shows the various implemented options to depict the converter operation in the frequency domain, namely the converter admittance expressed in Park and positive sequence coordinates (in Fig. 7(a) and Fig. 7(b), respectively), as well as the converter passivity index (in Fig. 7(c)). The results of the frequency domain analysis are consistent with the eigenvalue analysis, demonstrating that the admittance angle of the converter is outside of the (-$\degree,90$$\degree$) range in both coordinate systems and for the lowest value of the controller time constant. Additionally, the frequency region where this occurs coincides with the frequency of the unstable modes of Fig. 6(c). Finally, the passivity plot of Fig. 7(c) shows that the converter is non-passive in the same region, implying adverse interactions with the rest of the network in this frequency range and potential instability. The above frequency domain analysis features allow the seamless incorporation of black-box models to the stability analysis, provided that their respective frequency response is given to the software as an input, as explained in Section IV-D.
@@ VI-B INELFE System
This case study aims to showcase the capability of STAMP to perform analysis on AC/DC hybrid power systems and is based on the real-world HVDC interconnection between Spain and France, termed the INELFE link [francos2012inelfe]. During the $16^{\text{th}}$ of June 2015, the link suffered a tripping caused from a harmonic instability event, which this case study aims to reproduce [saad2016performance, saad2017resonances].
[…]
@@ VI-C Case Study Computation Time
Power flow
calculation
0.081 s 0.167 s 0.216 s 0.334 s
Subsystems linear
model calculation
0.441 s 0.277 s 2.044 s 6.242 s
Linear model
interconnection
0.009 s 0.008 s 0.36 s 2.528 s
Eigenvalue
calculation
0.1 s 0.109 s 0.616 s 3.895 s
@@
In this Section, the computation time that STAMP requires its operations is presented. These operations include the pre-calculation of the power flow for the linearization point evaluation, the calculation of the state space matrices of each individual linear subsystem, their interconnection and the final calculation of the eigenvalues from the complete system state matrix. The selected systems include the ones of Section VI-A and VI-B, as well as of two systems of larger size in order to demonstrate the scalability of the toolbox. The two larger models are a variation of the well-known New England benchmark, where 35% of the SGs are substituted with IBRs, as well as a reduced version of the Midwest benchmark system. Both the static and dynamic data that were used for the calculations can be found in [STAMP_Github]. All calculations were performed in a desktop computer, equipped with an AMD Ryzen Threadripper 2950X 16-core processor, operating at 3.5 GHz and a RAM of 32 GB.
@@ VII Conclusion
Acknowledgements
This work was partially funded by the REFORMING Project (PID2021-127788OA-I00) supported by Fondo Europeo de Desarrollo Regional/Ministerio de Ciencia e Innovación-Agencia Estatal de Investigación. The work of M. Cheah-Mañe and E. Prieto-Araujo is partially funded by the Serra Hunter Programme. O. Gomis-Bellmunt is an ICREA Academic Researcher.